OpenStudy (anonymous):

Convert to polar coordinates: x^2-3y+y^2=0

OpenStudy (psymon):

We just need to remember our conversions: \[r ^{2} = x ^{2} + y ^{2}\] \[x = rcos \theta \] \[y = rsin \theta \] Using those 3 conversions, think you can find out what to do from there?

OpenStudy (anonymous):

I think so. r^2 would be set equal to 3y so r=sqrt(3y), correct?

OpenStudy (anonymous):

Ok that still doesn't seem right, help out with next step? I can see that is is a circle, and that only r needs to be solved for. I'm not seeing the next step

OpenStudy (blockcolder):

No x's and y's should be left in the end, and there's still a y in r=sqrt(3y).

OpenStudy (dumbcow):

you need r in terms of theta \[r^{2}-3r \sin \theta = 0\] \[r = 3 \sin \theta\]